OpenStudy (anonymous):

What is the value of f[g(5)] for the functions f(x) = 2x + 1 and g(x) = 2x - 5. please help i have no idea what this means!!

OpenStudy (ash2326):

We have \[ f(x)=2x+1\] and \[g(x)= 2x-5\] We need to find f[g(5)] Let's find f[g(x)] first \[f(x)= 2x+1\] to find f[g(x)] substitute x = g(x) in f(x) we get \[f[g(x)]= 2g(x)+1\] now g(x)= 2x-5 so substituting value of g(x) \[f[g(x)]= 2(2x-5)+1\] Let's simplify \[f[g(x)]=4x-10+1=4x-9\] so \[f[g(x)]=4x-9\] We need to find f[g(5)], so substitute x=5 in f[g(x)] \[f[g(5)]= 4 (5)-9\] Can you solve the rest???

OpenStudy (anonymous):

I have no idea what you just did.....i just need the answer....but when i get the answer it comes out two.

OpenStudy (ash2326):

What didn't you get??